OpenStudy (anonymous):

g(x)=-x^2+2x, Find g(1/g(a))

OpenStudy (michele_laino):

please compute g(a) first

OpenStudy (michele_laino):

in order to that, it is suffice to set x=a into the formula for g(x)

OpenStudy (anonymous):

okay so then you get g(a)=a^2+2a @Michele_Laino

OpenStudy (michele_laino):

please, be careful with sign, since you have: \[g(a)=-a ^{2}+2a\]

OpenStudy (michele_laino):

Now, please you have to set \[x=\frac{ 1 }{ -2a ^{2}+2a }\] into your formula for g(x)

OpenStudy (michele_laino):

oops.. \[x=\frac{ 1 }{ -a ^{2}+2a }\]

OpenStudy (michele_laino):

what do you get?

OpenStudy (anonymous):

wouldnt that be the answer? I cant see anything I have to simplify

OpenStudy (michele_laino):

note that you have to compute this quantity: \[-\left( \frac{ 1 }{ -a ^{2}+2a } \right)^{2}+2\left( \frac{ 1 }{ -a ^{2}+2a } \right)\]

OpenStudy (michele_laino):

hint: \[-\left( \frac{ 1 }{ -a ^{2}+2a } \right)^{2}+\frac{ 2 }{ -a ^{2}+2a }=\]

OpenStudy (michele_laino):

=\[-\frac{ 1 }{ (-a ^{2}+2a )^{2}}+\frac{ 2 }{ (-a ^{2}+2a) }=\]

OpenStudy (michele_laino):

\[=\frac{ -1+2(-a ^{2}+2a) }{ (-a ^{2}+2a)^{2} }=...\] please continue